# Composable Interventions for Language Models

### v1

@@ 4.2 Composing Model Compression with Machine Unlearning
We next compose three model compression and three machine unlearning interventions. We measure unlearning success with WMDP (Section 3.1).
Compression hinders unlearning. Our main results in Figure 3 show that applying unlearning to a compressed model can degrade performance. With RMU, pruning the model before unlearning can significantly regress unlearning performance, especially at higher sparsity levels. We observe a similar trend with quantization when composing RMU and GPTQ. With GD, we similarly find that pruning after unlearning is best, though quantizing first can be optimal. We observe the opposite trend with GA, though GA’s poor unlearning performance, as seen in other studies [39, 60, 61], is a possible confounder. We overall find that unlearning a compressed model is more challenging but that the degree to which unlearning after compression hurts performance is highly contingent on the composition and compression level.
Order Sensitivity can determine overall composability. Table 4.3 reports the composability metrics of each unlearning technique. As measured by Order-free Error on WMDP, GD and RMU outperform GA, which struggles to unlearn without significantly degrading utility. GD exhibits higher WMDP Order Sensitivity for pruning than quantization, whereas RMU has low Order Sensitivity across all techniques except when quantized by GPTQ. Despite both GD and RMU’s successful unlearning, RMU has lower overall Order Sensitivity than GD, suggesting that while task performance is critical, it alone does not fully capture the practical challenges of applying interventions—RMU is more robust to order.
@@ 4.3 Composing Knowledge Editing with Machine Unlearning
4.4 General Trends in Composability
@@
Compression consistently hinders other interventions. Compressing an edited model can regress editing performance. Applying unlearning to a compressed model is often more challenging than unlearning the base model. These trends suggest that compression may modify how models encode knowledge in a way that makes targeted updates more challenging. Previous works have studied how compression can cause models to forget some knowledge [62, 63, 64, 65, 66]. Others have studied how editing modifies model internals in a way that makes the model more fragile [67, 68]. However, the relationships between how models are compressed and how knowledge is modified remain underexplored. An improved understanding of these dynamics may help guide more composable interventions and shed further light on how LMs encode knowledge.
RMU is the most composable unlearning technique. RMU has comparable unlearning performance to GD without regressing overall utility as harshly and while being more order invariant. A possible factor is the RMU training objective, which directs activations for inputs related to the unrelating to a random direction in activation space while minimally modifying unrelated activations. RMU achieves strong performance, even when only modifying the LM’s third layer. These results suggest that intervening on model representations [69] is a promising direction for effective and composable interventions.
MMLU is insufficient for measuring composability. MMLU is a common measure of a model’s general performance. However, similar scores on MMLU can mask significant differences in intervention performance. Across the board, interventions appear far more composable on MMLU than other metrics. We therefore advocate for thorough multi-metric and multi-dataset evaluations for composable interventions.
@@ 5 Conclusion
We conduct an extensive study of composing test-time interventions for language models. Such interventions are an important step towards keeping expensively trained models up-to-date over extended deployments with respect to factuality, efficiency, and social responsibility. Our study lays the foundation for a new range of intervention methods that are composable with one another, proposing two metrics for measuring the composability of interventions. Our setup aligns with the practical need to update the same models multiple times as use cases evolve and training remains expensive. We deploy this new framework to investigate composability between popular interventions spanning knowledge editing, machine unlearning, and model compression. We show that these interventions interact in non-trivial ways, pointing to concrete directions for future work. We further discover guidance for current practice: model compression harms other interventions, unlearning compressed models is especially hard, and the order in which interventions are applied changes their outcomes substantially. Our composability metrics also pave the way for formal investigations and practical principles. Finally, our framework is general and naturally extends to a broad range of test-time interventions.
@@ Appendix A Limitations
While our work provides the first look at how state-of-the-art methods compose, there are far more possible combinations of interventions, methods, models, and datasets than we can consider in one work. So some popular methods are omitted. We also prioritize broad experiments with many different intervention methods. We hope future works will closely examine individual methods to study, improve, and extend beyond the mechanisms that drive their composability. Lastly, we only experiment with Llama3-8B. While our code is extendable and Llama3-8B is popular, it is unknown whether our findings generalize between language models. We hope our codebase empowers others to explore more model sizes and architectures in future research.
@@ Appendix B Background on Interventions
Knowledge Editing. Knowledge editors aim to address factuality decay in LLMs: As the world changes, some facts an LLM learned during training become inaccurate [17, 70]. For example, a model trained before 2020 would still predict The latest pandemic in the US is “Swine Flu” until updated to predict “COVID”. There are four general approaches to model editing: 1) Memory-based methods that cache knowledge [71, 18], 2) Locate-then-Edit methods that selectively finetune parameter subsets [19, 20], 3) Hypernetwork methods that predict new model weights [72, 73, 22], and 4) Prompt editors that add facts to prompts [74, 75]. Most editors update singular facts [73, 76, 72, 19], though recent works apply many simultaneously [20, 71, 22]. Others works embrace the practical need for sequential edits [77, 18]. Recent works have also begun investigating unintended impacts of editing on pre-trained models [78, 79, 80, 81, 82, 83], while others are exploring multi-hop editing [84, 74, 75]. However, these works study impacts on model performance, not interactions with other types of interventions.
@@
Machine Unlearning. Machine Unlearning (MU) refers to a broad set of techniques for modifying a model post-training to remove the influence of specific training examples or other undesirable properties444Most MU techniques for LMs are considered approximate unlearning in that they aim to modify the model to approximate the behavior of it not being trained on the unlearning target without retraining from scratch. [53, 85, 27]. These undesirable properties are commonly referred to as the unlearning target. Crucially, MU must excel at causing the model to forget the unlearning target without adversely affecting overall utility by reducing performance on other domains [78]. Another requirement is that MU must cause models to deeply forget knowledge such that the unlearning target can not be elicited by adversaries or probing [28]. With the rise of generative models, recent work has explored MU for forgetting potentially copyrighted work [29], information about individuals [39], and potentially dangerous knowledge [86, 87, 40]. Existing works have studied which types of examples models forget when compressed [62], whether compression hinders unlearning [88], and whether compression in itself can be used as an unlearning technique [88]. While related to knowledge editing, the two areas have largely developed independently.
@@
Model Compression. To mitigate the huge computational and memory demands of LLMs, plenty of classical techniques in model compression have been explored, such as quantization [15, 37, 89], network pruning [16, 36, 90, 91], knowledge distillation [92, 93, 94], and low-rank factorization [34, 95, 96]. We focus on quantization and network pruning, which both make in-place updates to pre-trained model weights. Quantization involves lowering the bit-precision of model parameters, effectively shrinking model size and expediting inference. Early works, such as ZeroQuant [97] and LLM.int8 [98], focused on fine-tuning quantization granularity, showing promising results predominantly at higher bit-widths, like 8-bit. More recent innovations like GPTQ [15] and AWQ [37] have successfully pushed the boundaries by reducing bit-width to as low as 3 or 4 bits per weight, with negligible performance degradation. Pruning is another compression methods, which aims to eliminate redundant components like weights and channels. Techniques like SparseGPT [16] and Wanda [36] successfully maintain much of an LLM’s performance, even when discarding approximately 50% of its weights. OWL [90] achieves higher levels of sparsity by reallocating sparsities across layers.
@@ Appendix C Implementation Details
In this section we outline the set up and hyperparameters used for knowledge editing, unlearning, and compression.
@@ C.1.1 Knowledge Editing Metrics
Edit Success: Assesses the post-edit model’s ability to provide the correct, expected answer. An issue arises here with the different levels of verbosity for the models. An unedited model might return the correct token, but use a differently phrased answer or arrange the correct tokens differently. We therefore adopt and F1-style measure used in recent work [18] for this.
@@ C.1.2 Knowledge Editing Dataset
ZsRE [35]: This dataset focuses on context-free question-answering. Given a question related to a subject and a relation, the model’s task is to provide the correct object as the answer. [99] follow the procedure outlined in [44], to improve the locality sets used. As the initial locality sets were solely based on Natural Question annotations, we use the locality sets provided by [99].
@@ C.2.1 MEMIT Details
We use the state-of-the-art MEMIT [20] model editor, which applies batches of edits simultaneously. The editing process was applied to layers 4 through 8 of the model, with a clamp normalization factor set at 4. The learning parameters adhered closely to the original implementation: v_num_grad_steps was set to 25, accompanied by a learning rate (lr) of 0.5, and using the last layer for loss calculation. Additionally, a weight decay (weight_decay) of 0.001 was employed. The KL divergence contribution to the overall loss was controlled by a of 0.0625. Moreover, a second momentum adjustment was enabled, with an update weight of , to fine-tune the optimization process. The model generated a maximum length of 40 tokens and a batch size of 50, matching the number of edits being made. 10 repeats were made for each edit and the results averaged.
@@ C.3.2 Compressor details
SparseGPT [16]: an efficient one-shot pruning method tailored for large models. It converts the pruning process into solving large-scale sparse regression problems using an approximate solver. This approach enables rapid pruning on a single GPU with minimal accuracy loss, achieving 50-60% on large models.
Wanda [36]: is another popular method for pruning large language models that relies on a pruning metric that combines a weight’s magnitude with the norm of its corresponding input activations, determined from a small calibration dataset. The method focuses on selectively pruning weights within individual outputs of linear layers, aiming for high sparsity levels without modifying unpruned weights. Wanda is computationally efficient, executable in a single forward pass.
@@
GPTQ [15]: an algorithm designed for efficient weight quantization in large-scale models. It revises the weight quantization approach by quantizing weights in a fixed order rather than a greedy order, which shows minimal performance difference, especially in larger models. GPTQ introduced a novel method where each weight is quantized column-by-column, reducing computational complexity.
AWQ [37]: is based on the premise that not all weights are equally critical for model performance, and it identifies a small fraction of salient weights whose quantization significantly impacts model accuracy. This identification is done by analyzing activation distributions rather than weight distributions, under the rationale that weights linked to larger activation magnitudes are more crucial.
@@ C.4.1 Representation Misdirection Unlearning
Introduced in [40], RMU intervenes on a specific layer of the LM to perturb activations for inputs related to the unlearning target. This technique has achieved impressive performance on the WMDP dataset, with performance on the unlearning targets dropping to near-random while overall model utility (MMLU) minimally regresses. Wikitext [100] acts as the retain set.
RMU relies on multiple hyperparameters for selecting which layer to modify and how the loss should be calculated. We found that RMU is quite sensitive to hyperparameter choice, with most combinations either leaving the model unaffected or significantly harming model utility. Table 5 details the results from our hyperparameters search. We otherwise use the RMU repo’s555https://github.com/centerforaisafety/wmdp default settings.
@@ C.4.2 Gradient Ascent
GA, a simple and widely studied unlearning method [101, 61, 39, 102, 60], compels the model to minimize the likelihood of the correct token. The underlying intuition is that models will learn to avoid generating tokens associated with the unlearning target. We only train the final 16 layers of the model due to memory constraints.
@@ C.4.3 Gradient Difference
GD [39] aims to address this shortcoming by adding a retain loss term to the loss function. This term aims to maximize the likelihood of the correct token on an unrelated retain dataset. We use Wikitetx [100] as the retain set. As with GA, we only train the final 16 layers of the model.
Table 7 reports our hyperparameter search. Unlike [39], which found that GD provided only a marginal improvement over GA, we found that GD can perform significantly better if we upweight the retain loss term. Upweighting is necessary since it is easy for the model to significantly minimize the correct token likelihood, which causes it to drown out the retain loss term. Assigning a weight of 400 to the retain term solved this issue. This hyperparameter combination can make GD competitive with RMU, though still less composable (Section 4).

### v2

@@ 4.2 Composing Model Compression with Machine Unlearning
Compression hinders unlearning. Our main results in Figure 3 show that applying unlearning to a compressed model can degrade unlearning performance, measured via WMDP (Section 3.1). With RMU, pruning the model before unlearning can significantly regress unlearning performance, especially at higher sparsity levels. There is a similar trend with quantization when composing RMU and GPTQ. With GD, we similarly find that pruning after unlearning is best, though quantizing first can be best. We observe the opposite trend with GA, though GA’s poor unlearning performance, as seen in other studies (Maini et al., 2024; Zhang et al., 2024; Wang et al., 2024), is a possible confounder. Overall, unlearning a compressed model is hard, but the degree to which post-compression unlearning hurts performance is contingent on the composition and compression level.
Order Sensitivity can determine overall composability. Table 4.3 reports the composability metrics of each unlearning technique. Both GD and RMU’s successful unlearning (low Order-free Error), with RMU having lower overall Order Sensitivity than GD. While task performance is critical, it alone does not fully capture the practical challenges of applying interventions—RMU is more robust to order.
@@ 4.3 Composing Knowledge Editing with Machine Unlearning
4.4 Trends and Generalization Across Models
@@
Ablation experiments with Mistral 7B Instruct and Yi 1.5 9B Chat show that composition metrics are generally consistent across LMs. The exception being EditingCompression, as expected since performance has been shown to vary between models (Meng et al., 2023). EditingUnlearning and CompressionUnlearning consistently exhibit low Order Sensitivity. These results (detailed in Appendix E.1) suggest intervention composability trends often generalize across similar models. We also experiments across model families and sizes in Appendix E.2.
Compression consistently hinders other interventions. Compressing an edited model can regress editing performance. Unlearning on a compressed model is often more challenging than on the base model. The same trends appear with triplet interventions, Appendix E.4.1. These trends suggest compression may alter how models encode knowledge, making targeted updates harder. Previous works have explored how compression leads to knowledge loss (Hooker et al., 2019; Du et al., 2021; Azeemi et al., 2023; Hoang et al., 2023) and how editing modifies model internals, increasing fragility (Brown et al., 2023a; Gu et al., 2024b). However, the relationship between how models are compressed and how knowledge is modified remains underexplored. Understanding these dynamics could guide more composable interventions and illuminate how LMs encode knowledge.
MMLU is insufficient for measuring composability. MMLU is a common measure of a model’s general performance, but similar scores can mask significant differences in intervention performance. Interventions generally appear more composable on MMLU than on other metrics. Thus, we advocate for thorough multi-metric and multi-dataset evaluations for composable interventions.
@@ 5 Conclusion
We conduct an extensive study of composing test-time interventions for language models. Our setup aligns with the practical need to update the same models multiple times as use cases evolve and training remains expensive. We deploy this new framework to investigate composability between popular interventions spanning knowledge editing, machine unlearning, and model compression. We show that these interventions interact in non-trivial ways, pointing to concrete directions for future work. We further discover guidance for current practice: model compression harms other interventions, unlearning compressed models is especially hard, and the order in which interventions are applied changes their outcomes substantially. Our composability metrics also pave the way for formal investigations and practical principles. Finally, our framework is general and naturally extends to a broad range of test-time interventions.
@@ Appendix A Limitations
While our work provides the first look at how state-of-the-art methods compose, there are far more possible combinations of interventions, methods, models, and datasets than we can consider in one work. So some popular methods are omitted. We also prioritize broad experiments with many different intervention methods. We hope future works will closely examine individual methods to study, improve, and extend beyond the mechanisms that drive their composability. While we study how composability can vary across similar LMs (Section E.1), it is unknown whether our results generalize across model scale. We hope our codebase empowers others to explore more model sizes and architectures in future research.
@@ Appendix B Background on Interventions
Knowledge Editing. Knowledge editors aim to address factuality decay in LLMs: As the world changes, some facts an LLM learned during training become inaccurate (Mazzia et al., 2023; Wang et al., 2023c). For example, a model trained before 2020 would still predict The latest pandemic in the US is “Swine Flu” until updated to predict “COVID”. There are four general approaches to model editing: 1) Memory-based methods that cache knowledge (Mitchell et al., 2022b; Hartvigsen et al., 2023), 2) Locate-then-Edit methods that selectively finetune parameter subsets (Meng et al., 2022; 2023), 3) Hypernetwork methods that predict new model weights (Sinitsin et al., 2020; Mitchell et al., 2022a; Tan et al., 2024), and 4) Prompt editors that add facts to prompts (Zhong et al., 2023b; Cohen et al., 2024). Most editors update singular facts (Mitchell et al., 2022a; De Cao et al., 2021; Sinitsin et al., 2020; Meng et al., 2022), though recent works apply many simultaneously (Meng et al., 2023; Mitchell et al., 2022b; Tan et al., 2024). Others works embrace the practical need for sequential edits (Huang et al., 2023; Hartvigsen et al., 2023). Recent works have also begun investigating unintended impacts of editing on pre-trained models (Gu et al., 2024a; Hoelscher-Obermaier et al., 2023; Li et al., 2023; Hazra et al., 2024; Brown et al., 2023b; Hase et al., 2023), while others are exploring multi-hop editing (Powell et al., 2024; Zhong et al., 2023b; Cohen et al., 2024). However, these works study impacts on model performance, not interactions with other types of interventions.
@@
Machine Unlearning. Machine Unlearning (MU) refers to a broad set of techniques for modifying a model post-training to remove the influence of specific training examples or other undesirable properties444Most MU techniques for LMs are considered approximate unlearning in that they aim to modify the model to approximate the behavior of it not being trained on the unlearning target without retraining from scratch. (Cao & Yang, 2015; Shaik et al., 2023; Liu et al., 2024a). These undesirable properties are commonly referred to as the unlearning target. Crucially, MU must excel at causing the model to forget the unlearning target without adversely affecting overall utility by reducing performance on other domains (Gu et al., 2024c). Another requirement is that MU must cause models to deeply forget knowledge such that the unlearning target can not be elicited by adversaries or probing (Lynch et al., 2024). With the rise of generative models, recent work has explored MU for forgetting potentially copyrighted work (Eldan & Russinovich, 2023), information about individuals (Maini et al., 2024), and potentially dangerous knowledge (Zhao et al., 2023a; Liu et al., 2024b; Li et al., 2024a). Existing works have studied which types of examples models forget when compressed (Hooker et al., 2019), whether compression hinders unlearning (Jia et al., 2023), and whether compression in itself can be used as an unlearning technique (Jia et al., 2023). While related to knowledge editing, the two areas have largely developed independently.
@@
Model Compression. To mitigate the huge computational and memory demands of LLMs, plenty of classical techniques in model compression have been explored, such as quantization (Frantar et al., 2023; Lin et al., 2023; Xiao et al., 2023), network pruning (Frantar & Alistarh, 2023; Sun et al., 2023; Yin et al., 2023; Ma et al., 2023), knowledge distillation (Wang et al., 2023a; Saha et al., 2023; Hsieh et al., 2023), and low-rank factorization (hu2021lora; Xu et al., 2023a; Sharma et al., 2023). We focus on quantization and network pruning, which both make in-place updates to pre-trained model weights. Quantization involves lowering the bit-precision of model parameters, effectively shrinking model size and expediting inference. Early works, such as ZeroQuant (Yao et al., 2022) and LLM.int8 (Dettmers et al., 2022), focused on fine-tuning quantization granularity, showing promising results predominantly at higher bit-widths, like 8-bit. More recent innovations like GPTQ (Frantar et al., 2023) and AWQ (Lin et al., 2023) have successfully pushed the boundaries by reducing bit-width to as low as 3 or 4 bits per weight, with negligible performance degradation. Pruning is another compression methods, which aims to eliminate redundant components like weights and channels. Techniques like SparseGPT (Frantar & Alistarh, 2023) and Wanda (Sun et al., 2023) successfully maintain much of an LLM’s performance, even when discarding approximately 50% of its weights. OWL (Yin et al., 2023) achieves higher levels of sparsity by reallocating sparsities across layers.
@@ Appendix C Implementation Details
In this section we outline the set up and hyperparameters used for knowledge editing and compression: pruning and quantization.
@@ C.1.1 Knowledge Editing Metrics
Edit Success: Assesses the post-edit model’s ability to provide the correct, expected answer. An issue arises here with the different levels of verbosity for the models. An unedited model might return the correct token, but use a differently phrased answer or arrange the correct tokens differently. We therefore adopt and F1-style measure used in recent work (Hartvigsen et al., 2023) for this.
@@ C.1.2 Knowledge Editing Dataset
ZsRE (Levy et al., 2017): This dataset focuses on context-free question-answering. Given a question related to a subject and a relation, the model’s task is to provide the correct object as the answer. (Wang et al., 2023b) follow the procedure outlined in (Yao et al., 2023), to improve the locality sets used. As the initial locality sets were solely based on Natural Question annotations, we use the locality sets provided by (Wang et al., 2023b).
@@ C.2.1 MEMIT Details
We use the state-of-the-art MEMIT (Meng et al., 2023) model editor, which applies batches of edits simultaneously. The editing process was applied to layers 4 through 8 of the model, with a clamp normalization factor set at 4. The learning parameters adhered closely to the original implementation: v_num_grad_steps was set to 25, accompanied by a learning rate (lr) of 0.5, and using the last layer for loss calculation. Additionally, a weight decay (weight_decay) of 0.001 was employed. The KL divergence contribution to the overall loss was controlled by a of 0.0625. Moreover, a second momentum adjustment was enabled, with an update weight of , to fine-tune the optimization process. The model generated a maximum length of 40 tokens and a batch size of 50, matching the number of edits being made. 10 repeats were made for each edit and the results averaged.
@@ C.3.2 Compressor details
SparseGPT (Frantar & Alistarh, 2023): an efficient one-shot pruning method tailored for large models. It converts the pruning process into solving large-scale sparse regression problems using an approximate solver. This approach enables rapid pruning on a single GPU with minimal accuracy loss, achieving 50-60% on large models.
Wanda (Sun et al., 2023): is another popular method for pruning large language models that relies on a pruning metric that combines a weight’s magnitude with the norm of its corresponding input activations, determined from a small calibration dataset. The method focuses on selectively pruning weights within individual outputs of linear layers, aiming for high sparsity levels without modifying unpruned weights. Wanda is computationally efficient, executable in a single forward pass.
@@
GPTQ (Frantar et al., 2023): an algorithm designed for efficient weight quantization in large-scale models. It revises the weight quantization approach by quantizing weights in a fixed order rather than a greedy order, which shows minimal performance difference, especially in larger models. GPTQ introduced a novel method where each weight is quantized column-by-column, reducing computational complexity.
AWQ (Lin et al., 2023): is based on the premise that not all weights are equally critical for model performance, and it identifies a small fraction of salient weights whose quantization significantly impacts model accuracy. This identification is done by analyzing activation distributions rather than weight distributions, under the rationale that weights linked to larger activation magnitudes are more crucial.
@@ C.4.1 Representation Misdirection Unlearning
Introduced in (Li et al., 2024a), RMU intervenes on a specific layer of the LM to perturb activations for inputs related to the unlearning target. This technique has achieved impressive performance on the WMDP dataset, with performance on the unlearning targets dropping to near-random while overall model utility (MMLU) minimally regresses. Wikitext (Merity et al., 2016) acts as the retain set.
RMU relies on multiple hyperparameters for selecting which layer to modify and how the loss should be calculated. We found that RMU is quite sensitive to hyperparameter choice, with most combinations either leaving the model unaffected or significantly harming model utility. Table 5 details the results from our hyperparameters search for Llama-3 8B. We otherwise use the RMU repo’s555https://github.com/centerforaisafety/wmdp default settings.
@@ C.4.2 Gradient Ascent
GA, a simple and widely studied unlearning method (Golatkar et al., 2019; Wang et al., 2024; Maini et al., 2024; Yao et al., 2024; Zhang et al., 2024), compels the model to minimize the likelihood of the correct token. The underlying intuition is that models will learn to avoid generating tokens associated with the unlearning target. We only train the final 16 layers of the model due to memory constraints.
@@ C.4.3 Gradient Difference
GD (Maini et al., 2024) aims to address this shortcoming by adding a retain loss term to the loss function. This term aims to maximize the likelihood of the correct token on an unrelated retain dataset. We use Wikitext (Merity et al., 2016) as the retain set. As with GA, we only train the final 16 layers of the model.
Table 7 reports our hyperparameter search. Unlike (Maini et al., 2024), which found that GD provided only a marginal improvement over GA, we found that GD can perform significantly better if we upweight the retain loss term. Upweighting is necessary since it is easy for the model to significantly minimize the correct token likelihood, which causes it to drown out the retain loss term. Assigning a weight of 400 to the retain term solved this issue. This hyperparameter combination can make GD competitive with RMU, though still less composable (Section 4).